Please help!
1.) A blimp that usually goes y mph in still air, when a 10 mph tail wind is blowing?
2.) A boat traveling upstream on Pine Creek, which has a current of 2 mph. The boat usually goes z mph in still water.
3.) The plane flew to Omaha and then returned. The speed of the plane in still air was 100 mph more than the wind speed.
r t d

x + y 2 2(x + y)
x - y 3 3(x - y)
4.) The distance traveled by a jet in 12 hours at x mph. (symbolic expression)

Answer by ankor@dixie-net.com(12692)

1.) A blimp that usually goes y mph in still air, when a 10 mph tail wind is blowing?
(y+10) = speed with the wind
:
2.) A boat traveling upstream on Pine Creek, which has a current of 2 mph.
The boat usually goes z mph in still water.
(z-2) = speed upstream
:
3.) The plane flew to Omaha and then returned. The speed of the plane in still air was 100 mph more than the wind speed.
:
s = speed in still air
(s-100) = wind speed
with the wind:
s + (s-100) =
2s - 100 = speed with the wind
:
Against the wind
s - (s-100) =
s - s + 100 =
100 mph = speed against the wind
:
:
4.) The distance traveled by a jet in 12 hours at x mph. (symbolic expression)
dist = time * speed
d = 12x